### v4

# On the Necessity of Collaboration for Online Model Selection with Decentralized Data

Abstract
We consider online model selection with decentralized data over $M$ clients, and study the necessity of collaboration among clients. Previous work proposed various federated algorithms without demonstrating their necessity, while we answer the question from a novel perspective of computational constraints. We prove lower bounds on the regret, and propose a federated algorithm and analyze the upper bound. Our results show (i) collaboration is unnecessary in the absence of computational constraints on clients; (ii) collaboration is necessary if the computational cost on each client is limited to $o(K)$, where $K$ is the number of candidate hypothesis spaces. We clarify the unnecessary nature of collaboration in previous federated algorithms for distributed online multi-kernel learning, and improve the regret bounds at a smaller computational and communication cost. Our algorithm relies on three new techniques including an improved Bernstein’s inequality for martingale, a federated online mirror descent framework, and decoupling model selection and prediction, which might be of independent interest.
1 Introduction
Model selection which is a fundamental problem for offline machine learning focuses on how to select a suitable hypothesis space for a machine learning algorithm [1, 2, 3]. Model selection for online machine learning is called online model selection (OMS), such as model selection for online supervised learning [4, 5, 6, 7], model selection for online active learning [8], and model selection for contextual bandits [9, 10, 11]. We consider model selection for online supervised learning. Let =$\mathcal{F}=\{\mathcal{F}_{1},\ldots,\mathcal{F}_{K}\}$ contain $K$ hypothesis spaces and $\ell(\cdot,\cdot)$ be a loss function. For a sequence of examples $\{({\bf x}_{t},y_{t})\}_{t=1,\ldots,T}$, we aim to adapt to the case that the optimal hypothesis space $\mathcal{F}_{i^{\ast}}\in\mathcal{F}$ is given by an oracle and we run an online learning algorithm in $\mathcal{F}_{i^{\ast}}$. OMS can be defined by minimizing the regret, i.e.,
$\min_{f_{1},\ldots,f_{T}}\left(\sum^{T}_{t=1}\ell(f_{t}({\bf x}_{t}),y_{t})-\min_{f\in\mathcal{F}_{i^{\ast}}}\sum^{T}_{t=1}\ell(f({\bf x}_{t}),y_{t})\right),$
where $f_{t}\in\cup^{K}_{i=1}\mathcal{F}_{i}$ is the hypothesis used by an OMS algorithm at the $t$-th round. The optimal value of the regret depends on the complexity of $\mathcal{F}_{i^{\ast}}$ [4, 9].
In this work, we consider online model selection with decentralized data (OMS-DecD) over $M$ clients, in which each client observes a sequence of examples $\left\{\left({\bf x}^{(j)}_{t},y^{(j)}_{t}\right)\right\}_{t=1,\ldots,T}$, $j=1,\ldots,M$, and but does not share personalized data with others. There is a central server that coordinates the clients by sharing personalized models or gradients [12, 13, 14]. OMS-DecD captures some real-world applications where the data may be collected by sensors on $M$ different remote devices or mobile phones [15, 16, 17], or a local device can not store all of data due to low storage and thus it is necessary to store the data on more local devices [18, 19]. OMS-DecD can be defined by minimizing the following regret,
$\min_{f^{(j)}_{t},t=1,\ldots,T,j=1,\ldots,M}$ $\left(\sum^{M}_{j=1}\sum^{T}_{t=1}\ell\left(f^{(j)}_{t}\left({\bf x}^{(j)}_{t}\right),y^{(j)}_{t}\right)-\min_{f\in\mathcal{F}_{i^{\ast}}}\sum^{M}_{j=1}\sum^{T}_{t=1}\ell\left(f\left({\bf x}^{(j)}_{t}\right),y^{(j)}_{t}\right)\right),$
where $f^{(j)}_{t}\in\cup^{K}_{i=1}\mathcal{F}_{i}$ is the hypothesis adopted by the $j$-th client at the $t$-th round. Solving OMS-DecD must achieve two goals: G1 minimizing the regret, and G2 providing privacy protection.
A trivial approach is to use a noncooperative algorithm that independently runs a copy of an OMS algorithm on the $M$ clients. It naturally provides strong privacy protection, that is, it achieves G2, but suffers a regret bound that increases linearly with $M$. It is unknown whether it achieves G1. Another approach is federated learning which is a framework of cooperative learning with privacy protection and is provably effective in stochastic convex optimization [13, 20, 21, 22]. It is natural to ask:
Question 1.
Whether collaboration is effective for OMS-DecD.
The question reveals the hardness of OMS-DecD and is helpful to understand the limitations of federated learning. Previous work studied a special instance of OMS-DecD called distributed online multi-kernel learning (OMKL) where $\mathcal{F}_{i}$ is a reproducing kernel Hilbert space (RKHS), and proposed three federated OMKL algorithms including vM-KOFL, eM-KOFL [23] and POF-MKL [24]. The three algorithms also suffer regret bounds that increase linearly with $M$, and thus can not answer the question. If $K=1$, then OMS-DecD is equivalent to distributed online learning [25, 17, 16]. A noncooperative algorithm that independently runs online gradient descent (OGD) on each client achieves the two goals simultaneously [16]. Collaboration is unnecessary in the case of $K=1$.
In summary, previous work can not answer the question well. On one hand, previous work can not answer the question in the case of $K>1$. On the other hand, in the case of $K=1$, previous work has answered the question only using the statistical property of algorithms, i.e., the worst-case regret, but omitted the computational property which is very important for real-world applications.
1.1 Main Results
In this paper, we will answer the question from a new perspective of computational constraints on the problem (Section 5.5). Our main results are as follows.
  1. (1)
    An upper bound on the regret.  We propose a federated algorithm, FOMD-OMS, and prove an upper bound on the regret (Theorem 3). Besides, if $\mathcal{F}_{1},...,\mathcal{F}_{K}$ are RKHSs, then our algorithm improves the regret bounds of FOP-MKL [24] and eM-KOFL [23] at a smaller computational and communication cost. Table 2 summarizes the results.
  2. (2)
    Lower bounds on the regret.  We separately prove a lower bounds on the regret of any (possibly cooperative) algorithm and any noncooperative algorithm (Theorem 4).
  3. (3)
    A new perspective of computational constraints for Question 1.  By the upper bound and lower bounds, we conclude that (i) collaboration is unnecessary when there are no computational constraints on clients, thereby generalizing the result for distributed online learning, i.e., $K=1$; (ii) collaboration is necessary if the computational cost on each client is limited to $o(K)$ where irrelevant parameters are omitted. Our results clarify the unnecessary nature of collaboration in previous federated algorithms for distributed OMKL. Table 1 gives several results.
Table 1: Comparison with noncooperative algorithm (NCO-OMS). NCO-OMS independently runs a copy of an OMS algorithm on $M$ clients. $\Xi_{i^{\ast}}=\mathfrak{C}_{i^{\ast}}M\sqrt{T\ln{K}}$. $\mathfrak{C}_{i}$ measures the complexity of $\mathcal{F}_{i}$. $\mathfrak{C}_{i^{\ast}}$ measures the complexity of $\mathcal{F}_{i^{\ast}}$. $\mathfrak{C}=\max_{i\in\{1,\ldots,K\}}\mathfrak{C}_{i}\geq\mathfrak{C}_{i^{\ast}}$. The communication cost is the upload cost or download cost (bits). Comp-cost represents the per-round time complexity (s).
Constraint Algorithm Regret bound Comp-cost Communication cost
No(i.e., $R=T$) NCO-OMS $O\left(\mathfrak{C}_{i^{\ast}}M\sqrt{T\ln{(KT)}}\right)$ $O(K)$ 00
FOMD-OMS $O\left(\mathfrak{C}_{i^{\ast}}M\sqrt{T\ln{(KT)}}\right)$ $O(K)$ $O(KM)$
NCO-OMS $\tilde{O}\left(\sqrt{K}\left(\Xi_{i^{\ast}}+M\sqrt{\mathfrak{C}\mathfrak{C}_{i^{\ast}}T}\right)\right)$ $O(1)$ 00
FOMD-OMS $\tilde{O}\left(\Xi_{i^{\ast}}+\sqrt{MK}\sqrt{\mathfrak{C}\mathfrak{C}_{i^{\ast}}T}\right)$ $O(1)$ $O\left(M\log{K}\right)$
Table 2: Comparison with POF-MKL [24] and eM-KOFL [23]. $D$ is the number of random features [26]. $R$ is the rounds of communication. $\tilde{O}(\cdot)$ hides polylogarithmic factor in $T$. For the sake of simplicity, we omit the factor $O(\log{K})$ in the communication cost of eM-KOFL and FOMD-OMS. The unit of upload cost and download cost is bits.
Constraint Algorithm Regret bound Comp-cost Upload download
No ($R=T$) eM-KOFL $\tilde{O}\left(\mathfrak{C}M\sqrt{T\ln{K}}+\frac{\mathfrak{C}_{i^{\ast}}MT}{\sqrt{D}}\right)$ $O(DK)$ $O(DM)$ $O(DM)$
POF-MKL $\tilde{O}\left(\mathfrak{C}M\sqrt{KT}+\frac{\mathfrak{C}_{i^{\ast}}MT}{\sqrt{D}}\right)$ $O(DK)$ $O(DM)$ $O(DKM)$
FOMD-OMS $\tilde{O}\left(\Xi_{i^{\ast}}+\sqrt{\mathfrak{C}\mathfrak{C}_{i^{\ast}}MKT}+\frac{\mathfrak{C}_{i^{\ast}}MT}{\sqrt{D}}\right)$ $O(D)$ $O(DM)$ $O(DM)$
Yes ($R<T$) eM-KOFL - - - -
POF-MKL - - - -
FOMD-OMS $\tilde{O}\left(\frac{\Xi_{i^{\ast}}}{\sqrt{R/T}}+\frac{\sqrt{\mathfrak{C}\mathfrak{C}_{i^{\ast}}MK}T}{\sqrt{R}}+\frac{\mathfrak{C}_{i^{\ast}}MT}{\sqrt{D}}\right)$ $O(D)$ $O\left(\frac{DMR}{T}\right)$ $O\left(\frac{DMR}{T}\right)$
1.2 Technical Challenges
There are two main technical challenges on designing a federated online model selection algorithm.
The first challenge lies in obtaining high-probability regret bounds that adapt to the complexity of individual hypothesis space, a fundamental problem in online model selection [4]. While acquiring expected regret bounds that adapt to the complexity of individual hypothesis spaces is straightforward, the crux is to derive high-probability bounds from expected bounds. To this end, we introduce a new Bernstein’s inequality for martingale (Lemma 1), which might be of independent interest.
The second challenge involves achieving a per-round communication cost of $o(K)$. To tackle this challenge, we propose two techniques: (i) decoupling model selection and prediction; (ii) an algorithmic framework, named FOMD-No-LU, which might be of independent interest. Specifically, when clients execute model selection, server must broadcast an aggregated probability distribution, denoted by ${\bf p}\in\mathbb{R}^{K}$, to clients, naturally incurring a $O(K)$ download cost. Our algorithm conducts model selection on server and makes predictions on clients, thereby eliminating the need to broadcast the aggregated probability distribution to clients. Additionally, if we use the local updating approach [25, 16], then server must broadcast $K$ aggregated models to clients, also resulting in a $O(K)$ download cost [24]. By utilizing FOMD-No-LU, our algorithm only broadcasts the selected models to clients and can achieve a $o(K)$ download cost.
2 Related work
Previous work has studied the necessity of collaboration for distributed bandit convex optimization [16], where a federated algorithmic framework named FEDPOSGD was proposed. Although the regret bounds of FEDPOSGD are smaller than some noncooperative algorithms, there is not a lower bound on the regret of any noncooperative algorithm [16]. Moreover, the regret analysis of FEDPOSGD is based on the analysis for federated online gradient descent that is not applicable to our algorithm, FOMD-OMS. The regret analysis of FOMD-OMS requires the analysis for federated online mirror descent with negative entropy regularizer.
Our work is also different from federated bandits, such as federated $K$-armed bandits [27] and federated linear contextual bandits [28]. For OMS-DecD, we do not assume that the examples $({\bf x}^{(j)}_{t},y^{(j)}_{t})$, $t=1,\ldots,T$, on each client are independent and identically distributed (i.i.d.). In contrast, in both federated $K$-armed bandits or federated linear contextual bandits, the rewards must be i.i.d., thereby making collaboration effective. This is similar to the approach used in federated stochastic optimization. However, this may not hold true for OMS-DedD. Therefore, it is a distinctive problem for OMS-DecD to study whether collaboration is effective.
3 Preliminaries and Problem Setting
3.1 Notations
Let $\mathcal{X}=\{\mathbf{x}\in\mathbb{R}^{d}|\|\mathbf{x}\|_{2}<\infty\}$ be an instance space, $\mathcal{Y}=\{y\in\mathbb{R}:|y|<\infty\}$ be an output space, and $\mathcal{I}_{T}=\{(\mathbf{x}_{t},y_{t})\}_{t\in[T]}$ be a sequence of examples, where $[T]=\{1,\ldots,T\}$, $\mathbf{x}_{t}\in\mathcal{X}$ and $y_{t}\in\mathcal{Y}$. Let $S=\{s_{1},s_{2},\ldots\}$ be a finite set, $\mathrm{Uni}(S)$ be the uniform distribution over the elements in $S$ and $s_{[T]}$ be the abbreviation of the sequence $s_{1},s_{2},\ldots,s_{T}$. Denote by $\mathbb{P}[A]$ the probability that an event $A$ occurs, $a\wedge b=\min\{a,b\}$, $a\vee b=\max\{a,b\}$ and $\log(a)=\log_{2}(a)$. Let $\psi_{t}(\cdot):\Omega\rightarrow\mathbb{R},t\in[T]$ be a sequence of time-variant strongly convex regularizers defined on a domain $\Omega$. The Bregman divergence denoted by $\mathcal{D}_{\psi_{t}}(\cdot,\cdot)$, associated with $\psi_{t}(\cdot)$ is defined by
$\forall{\bf u},{\bf v}\in\Omega,\quad\mathcal{D}_{\psi_{t}}(\mathbf{u},\mathbf{v})=\psi_{t}(\mathbf{u})-\psi_{t}(\mathbf{v})-\langle\nabla\psi_{t}(\mathbf{v}),\mathbf{u}-\mathbf{v}\rangle.$
3.2 Online Model Selection (OMS)
Let =$\mathcal{F}=\{\mathcal{F}_{1},...,\mathcal{F}_{K}\}$ contain $K$ hypothesis spaces where
$\mathcal{F}_{i}=\left\{f({\bf x})={\bf w}^{\top}\phi_{i}({\bf x}):\phi_{i}({\bf x})\in\mathbb{R}^{d_{i}},\|{\bf w}\|_{2}\leq U_{i}\right\},$ (1)
where $\|\cdot\|_{2}$ is the $L_{2}$ norm. Let $\mathcal{F}_{i^{\ast}}\in\mathcal{F}$ be the optimal but unknown hypothesis space for a given $\mathcal{I}_{T}$. OMS can be defined as follows: generating a sequence of hypotheses $f_{[T]}$ that minimizes the following regret,
$\forall i\in[K],\quad\mathrm{Reg}(\mathcal{F}_{i})=\sum^{T}_{t=1}\ell(f_{t}({\bf x}_{t}),y_{t})-\min_{f\in\mathcal{F}_{i}}\sum^{T}_{t=1}\ell(f({\bf x}_{t}),y_{t}),$
where $f_{t}\in\cup^{K}_{i=1}\mathcal{F}_{i}$. The optimal hypothesis space $\mathcal{F}_{i^{\ast}}$ must contain a good hypothesis and has a low complexity [4, 9], and is defined by
$\mathcal{F}_{i^{\ast}}=\mathop{\arg\min}_{\mathcal{F}_{i}\in\mathcal{F}}\left[\min_{f\in\mathcal{F}_{i}}\sum^{T}_{t=1}\ell(f({\bf x}_{t}),y_{t})+\Theta\left(\sqrt{T\cdot\mathfrak{C}_{i}}\right)\right],$
where $\mathfrak{C}_{i}$ measures the complexity of $\mathcal{F}_{i}$, such as $U_{i}$ and $d_{i}$.
OMS is more challenge than online learning, since we not only learn the optimal hypothesis space, but also learn the optimal hypothesis in the space. Next we give some examples of OMS.
Example 1 (Online Hyper-parameters Tuning).
Let $\mathcal{F}_{i}$ consist of linear functions of the form
$\mathcal{F}_{i}=\left\{f({\bf x})=\langle{\bf w},{\bf x}\rangle,\|{\bf w}\|_{2}\leq U_{i}\right\},$
where $U_{i}>0$ is a regularization parameter. Let $\mathcal{U}=\{U_{i},i\in[K]:U_{1}<U_{2}<\ldots<U_{K}\}$. The hypothesis spaces are nested, i.e., $\mathcal{F}_{1}\subseteq\mathcal{F}_{2}\subseteq\ldots\subseteq\mathcal{F}_{K}.$ The optimal regularization parameter $U_{i^{\ast}}\in\mathcal{U}$ corresponds to the optimal hypothesis space $\mathcal{F}_{i^{\ast}}\in\mathcal{F}$.
Example 2 (Online Kernel Selection [29, 7]).
Let $\kappa_{i}(\cdot,\cdot):\mathbb{R}^{d}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$ be a positive semidefinite kernel function, and $\phi_{i}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d_{i}}$ be the associated feature mapping. $\mathcal{F}_{i}$ is the RKHS associated with $\kappa_{i}$, i.e.,
$\mathcal{F}_{i}=\left\{f({\bf x})=\langle{\bf w},\phi_{i}({\bf x})\rangle:\|{\bf w}\|_{2}\leq U_{i}\right\}.$
The optimal kernel function $\kappa_{i^{\ast}}\in\{\kappa_{1},\ldots,\kappa_{K}\}$ corresponds to the optimal RKHS $\mathcal{F}_{i^{\ast}}\in\mathcal{F}$.
Example 3 (Online Pre-trained Classifier Selection [8]).
Generally, $\mathcal{F}_{i}$ can be a well-trained machine learning model. Let $\mathcal{F}$ contain $K$ pre-trained classifiers. For a new instance ${\bf x}_{t}$, we select a (combinational) pre-trained classifier and make a prediction. The selection of a pre-trained classifier has an important implication in practical scenarios.
3.3 Online Model Selection with Decentralized Data (OMS-DecD)
We formally define OMS-DecD as follows. Assuming that there are $M$ clients and a server. At any round $t$, each client observes an instance ${\bf x}^{(j)}_{t}$, and selects a hypothesis $f^{(j)}_{t}\in\cup^{K}_{i=1}\mathcal{F}_{i}$, $j\in[M]$. Then clients output predictions $\{f^{(j)}_{t}({\bf x}^{(j)}_{t})\}^{M}_{j=1}$. The goal is to minimize the following regret
$\forall i\in[K],\quad\mathrm{Reg}_{D}(\mathcal{F}_{i})=\sum^{T}_{t=1}\sum^{M}_{j=1}\ell\left(f^{(j)}_{t}({\bf x}^{(j)}_{t}),y^{(j)}_{t}\right)-\min_{f\in\mathcal{F}_{i}}\sum^{T}_{t=1}\sum^{M}_{j=1}\ell\left(f({\bf x}^{(j)}_{t}),y^{(j)}_{t}\right),$
where $y^{(j)}_{t}$ is the label or true output. Each client can not share personalized data with others, but can share personalized models or gradients via the central server. For simplicity, we define OMS-DecD in Protocol 1.
Algorithm 1 Protocol 1 OMS-DecD
1:  for $t=1,2,\ldots,T$ do
2:     for $j=1,\ldots,M$ in parallel do
3:        The adversary sends ${\bf x}^{(j)}_{t}$ to the $j$-th client
4:        The learner selects a hypothesis space $\mathcal{F}_{I_{t}}\in\mathcal{F}$
5:        The learner selects $f^{(j)}_{t}\in\mathcal{F}_{I_{t}}$ and outputs $f^{(j)}_{t}({\bf x}^{(j)}_{t})$
6:        The learner observes the true output $y^{(j)}_{t}$
7:     end for
8:  end for
4 FOMD-No-LU
In this section, we propose a federated algorithmic framework, FOMD-No-LU (Federated Online Mirror Descent without Local Updating) for online collaboration.
4.1 Federated Algorithmic Framework
Let $\Omega$ be a convex and bounded decision set. At any round $t$, each client $j\in[M]$ first selects a decision ${\bf u}^{(j)}_{t}\in\Omega$, and then observes a loss function $l^{(j)}_{t}(\cdot):\Omega\rightarrow\mathbb{R}$. The client computes the loss $l^{(j)}_{t}({\bf u}^{(j)}_{t})$ and an estimator of the gradient denoted by $\tilde{g}^{(j)}_{t}$ (or the gradient denoted by $g^{(j)}_{t}$). To reduce the communication cost, we adopt the intermittent communication (IC) protocol [30], in which the clients communicate with the server every $N$ rounds. Assuming that $T=N\times R$ where $N,R\in\mathbb{Z}$, the IC protocol limits the rounds of communication to $R$.
We divide $[T]$ into $R$ disjoint sub-intervals denoted by $\{T_{r}\}^{R}_{r=1}$, in which
$T_{r}=\left\{(r-1)N+1,(r-1)N+2,\ldots,rN\right\}.$ (2)
For any $t\in T_{r}$, all clients always select the initial decision,
$\forall j\in[M],~{}\forall t\in T_{r},\quad{\bf u}^{(j)}_{t}={\bf u}^{(j)}_{(r-1)N+1}.$ (3)
At the end of the $rN$-round, all of clients send $\frac{1}{N}\sum_{t\in T_{r}}\tilde{g}^{(j)}_{t}$, $j\in[M]$ to server. Then the server updates the decision using online mirror descent framework [31, 32],
${\bf u}_{t}$ $=\frac{1}{M}\sum^{M}_{j=1}{\bf u}^{(j)}_{t},$ (4)
$\bar{g}_{t}$ $=\frac{1}{M}\sum^{M}_{j=1}\left(\frac{1}{N}\sum_{t\in T_{r}}\tilde{g}^{(j)}_{t}\right),$ (5)
$\nabla_{\bar{{\bf u}}_{t+1}}\psi_{t}(\bar{{\bf u}}_{t+1})$ $=\nabla_{{\bf u}_{t}}\psi_{t}({\bf u}_{t})-\bar{g}_{t},$ (6)
${\bf u}_{t+1}$ $=\mathop{\arg\min}_{{\bf u}\in\Omega}\mathcal{D}_{\psi_{t}}({\bf u},\bar{{\bf u}}_{t+1}).$ (7)
(4)-(6) is called model averaging [13] and shows the collaboration among clients. Finally, the server may broadcast ${\bf u}_{t+1}$ to all clients, i.e.,
$\forall j\in[M],\quad{\bf u}^{(j)}_{t+1}={\bf u}_{t+1}.$
Let the initial decision ${\bf u}^{(j)}_{1}={\bf u}_{1}$ for all $j\in[M]$, then it must be ${\bf u}^{(j)}_{t}={\bf u}_{t}$ for all $t\in[T]$. Thus (4) is unnecessary, and the clients do not transmit ${\bf u}^{(j)}_{t}$ to server. The pseudo-code of FOMD-No-LU is shown in Algorithm 2.
Algorithm 2 FOMD-No-LU
0:  $\Omega$.
0:  ${\bf u}^{(j)}_{1},j\in[M]$
1:  for $r=1,2,\ldots,R$ do
2:     for $t=(r-1)N+1,\ldots,rN$ do
3:        for $j=1,\ldots,M$ in parallel do
4:           Selecting ${\bf u}^{(j)}_{(r-1)N+1}$
5:           Observing loss function $l^{(j)}_{t}(\cdot)$
6:           Computing gradient (or an estimator of gradient) $\tilde{g}^{(j)}_{t}$
7:           if $t==rN$ then
8:              Transmitting $\frac{1}{N}\sum_{t\in T_{r}}\tilde{g}^{(j)}_{t}$ to server
9:           end if
10:        end for
11:        if $t==rN$ then
12:           Server computes ${\bf u}_{t+1}$ following (5), (6) and (7)
13:           Server may broadcast ${\bf u}_{t+1}$: ${\bf u}^{(j)}_{t+1}={\bf u}_{t+1}$, $j\in[M]$
14:        end if
15:     end for
16:  end for
4.2 Regret Bound
We give the regret bounds of FOMD-No-LU.
Theorem 1.
Let $R=T$. Assuming that $l^{(j)}_{t}:\Omega\rightarrow\mathbb{R},t\in[T],j\in[M]$ is convex. Let $g^{(j)}_{t}=\nabla_{{\bf u}^{(j)}_{t}}l^{(j)}_{t}({\bf u}^{(j)}_{t})$ and $\tilde{g}^{(j)}_{t}$ be an estimator of $g^{(j)}_{t}$. At any round $t$, let ${\bf q}_{t+1}$ and ${\bf r}_{t+1}$ be two auxiliary decisions defined as follows,
$\nabla_{{\bf q}_{t+1}}\psi_{t}({\bf q}_{t+1})=$ $\nabla_{{\bf u}_{t}}\psi_{t}({\bf u}_{t})-2\sum^{M}_{j=1}\frac{\tilde{g}^{(j)}_{t}-g^{(j)}_{t}}{M},$ (8)
$\nabla_{{\bf r}_{t+1}}\psi_{t}({\bf r}_{t+1})=$ $\nabla_{{\bf u}_{t}}\psi_{t}({\bf u}_{t})-\frac{2}{M}\sum^{M}_{j=1}g^{(j)}_{t}.$ (9)
Then FOMD-No-LU guarantees that,
$\forall{\bf v}\in\Omega,\quad\frac{1}{M}\sum^{T}_{t=1}\sum^{M}_{j=1}\left(l^{(j)}_{t}({\bf u}^{(j)}_{t})-l^{(j)}_{t}({\bf v})\right)$
$\leq$ $\underbrace{\sum^{T}_{t=1}\left[\mathcal{D}_{\psi_{t}}({\bf v},{\bf u}_{t})-\mathcal{D}_{\psi_{t}}({\bf v},{\bf u}_{t+1})+\frac{\mathcal{D}_{\psi_{t}}({\bf u}_{t},{\bf r}_{t+1})}{2}\right]}_{\Xi_{1}}+\underbrace{\sum^{T}_{t=1}\left[\frac{\mathcal{D}_{\psi_{t}}({\bf u}_{t},{\bf q}_{t+1})}{2}+\sum^{M}_{j=1}\frac{\left\langle\tilde{g}^{(j)}_{t}-g^{(j)}_{t},{\bf u}_{t}-{\bf v}\right\rangle}{M}\right]}_{\Xi_{2}}.$
It is intriguing that the regret bound comprises two components: the first part, $\Xi_{1}$, cannot be reduced by collaboration, while the second part, $\Xi_{2}$, highlights the benefits of collaboration. $\Xi_{1}$ is the regret induced by exact gradients, while $\Xi_{2}$ is the regret induced by estimated gradients and shows how collaboration controls the regret. It is worth mentioning that Theorem 1 gives a general regret bound, from which various types of regret bounds can be readily derived by instantiating the decision set $\Omega$ and the regularizer $\psi_{t}(\cdot)$. For instance, if $\Omega=\mathcal{F}_{i}$ where $\mathcal{F}_{i}$ follows Example 1, $\psi_{t}({\bf v})=\frac{1}{2\lambda}\|{\bf v}\|^{2}_{2}$ and $\mathbb{E}[\|\tilde{g}^{(j)}_{t}\|^{2}_{2}]\leq C\|g^{(j)}_{t}\|^{2}_{2}$, then FOMD-No-LU becomes a federated online descent descent. It is easy to give a $O(MU_{i}\sqrt{(1+\frac{C}{M})T})$ expected regret from Theorem 1. Besides, $N>1$ increases the regret and shows the trade-off between communication cost and regret bound.
Theorem 1 requires a novel analysis on how the bias of estimators, i.e., $\sum^{M}_{j=1}\|\tilde{g}^{(j)}_{t}-g^{(j)}_{t}\|^{2}_{2}$, is controlled by cooperation. To this end, we introduce two virtual decisions ${\bf q}_{t+1}$ and ${\bf r}_{t+1}$ that are updated by $2\sum^{M}_{j=1}\frac{\tilde{g}^{(j)}_{t}-g^{(j)}_{t}}{M}$ and $2\sum^{M}_{j=1}\frac{g^{(j)}_{t}}{M}$, respectively. Previous federated online mirror descent uses exact gradients $g^{(j)}_{t},j\in[M]$ [25]. Thus its analysis is different from ours.
Theorem 2.
Let $R<T$ and =$\mathcal{R}=\{N,2N,\ldots,RN\}$. At any round $t\in\mathcal{R}$, let ${\bf q}_{t+1}$ and ${\bf r}_{t+1}$ be two auxiliary decisions which follow (8) and (9). Under the assumptions in Theorem 1, FOMD-No-LU guarantees that,
$\forall{\bf v}\in\Omega,\quad\frac{1}{NM}\sum^{T}_{t=1}\sum^{M}_{j=1}\left(l^{(j)}_{t}({\bf u}^{(j)}_{t})-l^{(j)}_{t}({\bf v})\right)$
$\leq$ $\sum_{t\in\mathcal{R}}\left[\mathcal{D}_{\psi_{t}}({\bf v},{\bf u}_{t})-\mathcal{D}_{\psi_{t}}({\bf v},{\bf u}_{t+1})+\frac{\mathcal{D}_{\psi_{t}}({\bf u}_{t},{\bf r}_{t+1})}{2}\right]+\sum_{t\in\mathcal{R}}\left[\frac{\mathcal{D}_{\psi_{t}}({\bf u}_{t},{\bf q}_{t+1})}{2}+\sum^{M}_{j=1}\frac{\left\langle\tilde{g}^{(j)}_{t}-g^{(j)}_{t},{\bf u}_{t}-{\bf v}\right\rangle}{M}\right].$
It is obvious that $N>1$ increases the regret, that is, the reduction on the communication cost is at the cost of regret, which shows the trade-off between communication cost and regret bound. We will explicitly give the trade-off.
4.3 Comparison with Previous Work
In fact, FOMD-No-LU adopts the batching technique [33], that is, it divides $[T]$ into $R$ sub-intervals and executes (3) during each sub-intervals. The batching technique (also known as mini-batch) has been used in the multi-armed bandit problem [34] and distributed stochastic convex optimization [35, 36]. We use the batching technique for the first time to distributed online learning.
FOMD-No-LU is different from FedOMD (federated online mirror descent) [25]. (i) FedOMD only transmits exact gradients, while FOMD-No-LU can transmit estimators of gradient. Thus the regret bound of FedOMD did not contain $\Xi_{2}$ in Theorem 1. (ii) FedOMD uses local updating, such as local OGD [16] and local SGD [13, 22]. Thus FedOMD induces the client drift, i.e., ${\bf u}^{(j)}_{t}\neq{\bf u}_{t}$. Besides, if we use FedOMD, then the download cost is in $O(MK)$.
5 OMS-DecD without Communication Constraints
At a high level, our algorithm comprises two components both of which are critical for achieving a communication cost in $o(K)$: (i) decoupling model selection and online prediction; (ii) collaboratively updating decisions within the framework of FOMD-No-LU.
5.1 Decoupling Model Selection and Prediction
5.1.1 Model Selection on Server
At any round $t$, server maintains $K$ hypotheses $\{f^{(j)}_{t,i}\in\mathcal{F}_{i}\}^{K}_{i=1}$ and a probability distribution ${\bf p}^{(j)}_{t}$ over the $K$ hypotheses for all $j\in[M]$. The model selection process aims to select a hypothesis from $\{f^{(j)}_{t,i}\}^{K}_{i=1}$ and then predicts the output of ${\bf x}^{(j)}_{t}$. An intuitive idea is that, for each $j\in[M]$, the client samples a hypothesis following ${\bf p}^{(j)}_{t}$. However, such an approach requires that server broadcasts ${\bf p}^{(j)}_{t}$ to clients, and will cause a download cost in $O(K)$.
The sampling operation (or model selection process) can be executed on server. Specifically, server just broadcasts the selected hypotheses, and thus saves the communication cost. For each $j\in[M]$, server selects $J\in[2,K]$ hypotheses denoted by $f^{(j)}_{t,A_{t,a}}$, $a\in[J]$ where $A_{t,a}\in[K]$. For simplicity, let $O^{(j)}_{t}=\{A_{t,1},\ldots,A_{t,J}\}$. We instantiate ${\bf u}_{t}={\bf p}_{t}$ in FOMD-No-LU. Then FOMD-No-LU ensures ${\bf p}^{(j)}_{t}={\bf p}_{t}$ for all $j\in[M]$. We sample $A_{t,1},\ldots,A_{t,J}$ in order and follow (10).
$\begin{split}&A_{t,1}\sim{\bf p}_{t},\\ &A_{t,a}\sim\mathrm{Uni}([K]\setminus\{A_{t,1},\ldots,A_{t,a-1}\}),~{}a\in[2,J].\end{split}$ (10)
It is easy to prove that
$\forall i\in[K],\quad\mathbb{P}\left[i\in O^{(j)}_{t}\right]=\frac{K-J}{K-1}p_{t,i}+\frac{J-1}{K-1}.$
Server samples $O^{(j)}_{t}$ for all $j\in[M]$ and thus must independently execute (10) $M$ times which only pays an additional computational cost in $O(M\log{K})$. The factor $\log{K}$ arises from the process of sampling a number from $\{1,...,K\}$. Server only sends $f^{(j)}_{A_{t,a}}$, $a\in[J]$ to the $j$-th client. It is worth mentioning that server does not send ${\bf p}_{t}$. The total download cost is $O(\sum^{M}_{j=1}\sum^{J}_{a=1}(d_{A_{t,a}}+\log{K}))$. If $J$ is independent of $K$, then the download cost is only $O(M\log{K})$.
5.1.2 Prediction on Clients
For each $j\in[M]$, the $j$-th client receives $f^{(j)}_{A_{t,a}}$, $a\in[J]$, and uses $f^{(j)}_{t,A_{t,1}}$ to output a prediction, i.e.,
$\hat{y}^{(j)}_{t}=f^{(j)}_{t,A_{t,1}}\left({\bf x}^{(j)}_{t}\right)=\left\langle{\bf w}^{(j)}_{t,A_{t,1}},\phi_{A_{t,1}}({\bf x}^{(j)}_{t})\right\rangle,$
where we assume that $f^{(j)}_{t,i}$ is parameterized by ${\bf w}^{(j)}_{t,i}\in\mathbb{R}^{d_{i}}$ (see (1)). After observing the true output $y^{(j)}_{t}$, the client suffers a loss $\ell(f^{(j)}_{t,A_{t,1}}({\bf x}^{(j)}_{t}),y^{(j)}_{t})$.
It is worth mentioning that the other $J-1$ hypotheses $f^{(j)}_{t,A_{t,a}}$, $a\geq 2$ are just used to obtain more information on the loss function. We will explain more in the following subsection. Thus we do not cumulate the loss $\ell(f^{(j)}_{t,A_{t,a}}({\bf x}^{(j)}_{t}),y^{(j)}_{t})$, $a\geq 2$.
5.2 Online Collaboration Updating
We use FOMD-No-LU to update the sampling probabilities and the hypotheses.
5.2.1 Updating sampling probabilities
For each $j\in[M]$, let ${\bf c}^{(j)}_{t}=(c^{(j)}_{t,1},\ldots,c^{(j)}_{t,K})$ where $c^{(j)}_{t,i}=\ell(f^{(j)}_{t,i}({\bf x}^{(j)}_{t}),y^{(j)}_{t})$ is the loss of $f^{(j)}_{t,i}$, $i\in[K]$. The $j$-th client will send $c^{(j)}_{t,i}$, $i\in O^{(j)}_{t}$, to server. Since $c^{(j)}_{t,i}$, $i\notin O^{(j)}_{t}$ can not be observed, it is necessary to construct an estimated loss vector $\tilde{{\bf c}}^{(j)}_{t}=(\tilde{c}^{(j)}_{t,1},\ldots,\tilde{c}^{(j)}_{t,K})$ where
$\tilde{c}^{(j)}_{t,i}=\frac{c^{(j)}_{t,i}}{\mathbb{P}[i\in O^{(j)}_{t}]}\cdot\mathbb{I}_{i\in O^{(j)}_{t}},i\in[K].$
It is easy to prove that $\mathbb{E}_{t}\left[\tilde{c}^{(j)}_{t,i}\right]=c^{(j)}_{t,i}$ and $\mathbb{E}_{t}\left[(\tilde{c}^{(j)}_{t,i})^{2}\right]\leq\frac{K-1}{J-1}(c^{(j)}_{t,i})^{2}$ where $\mathbb{E}_{t}[\cdot]:=\mathbb{E}\left[\cdot|O^{(j)}_{[t-1]}\right]$. Thus sampling $A_{t,a}$, $a\geq 2$ reduces the variance of the estimators which is equivalent to obtain more information on the true loss.
Server aggregates $\tilde{{\bf c}}^{(j)}_{t}$, $j\in[M]$ and updates ${\bf p}_{t}$ following (5)-(7). Let $\Delta_{K}$ be the $(K-1)$-dimensional simplex, $\Omega=\Delta_{K}$ and $\tilde{g}^{(j)}_{t}=\tilde{{\bf c}}^{(j)}_{t}$. Then the server executes (11).
$\left\{\begin{split}\bar{\bf c}_{t}&=\frac{1}{M}\sum^{M}_{j=1}\tilde{{\bf c}}^{(j)}_{t},\\ \nabla_{\bar{{\bf p}}_{t+1}}\psi_{t}(\bar{{\bf p}}_{t+1})&=\nabla_{{\bf p}_{t}}\psi_{t}({\bf p}_{t})-\bar{{\bf c}}_{t},\\ {\bf p}_{t+1}&=\mathop{\arg\min}_{{\bf p}\in\Delta_{K}}\mathcal{D}_{\psi_{t}}({\bf p},\bar{{\bf p}}_{t+1}),\\ \psi_{t}(\mathbf{p})&=\sum^{K}_{i=1}\frac{C_{i}}{\eta_{t}}p_{i}\ln{p_{i}},\\ \end{split}\right.$ (11)
where $\psi_{t}(\mathbf{p})$ is the weighted negative entropy regularizer [37], $C_{i}>0$ is the weight and $\eta_{t}>0$ is a time-variant learning rate. $C_{i}$ satisfies that $\max_{t}c^{(j)}_{t,i}\leq C_{i}$ for all $j\in[M]$. The server does not broadcast ${\bf p}_{t+1}$.
5.2.2 Updating hypotheses
For each $j\in[M]$ and $i\in[K]$, let $\nabla^{(j)}_{t,i}=\nabla_{{\bf w}^{(j)}_{t,i}}\ell\left(\left\langle{\bf w}^{(j)}_{t,i},\phi_{i}({\bf x}^{(j)}_{t})\right\rangle,y^{(j)}_{t}\right)$. Since $\nabla^{(j)}_{t,i},i\notin O^{(j)}_{t}$ are unknown, it is necessary to construct an estimator of the gradient, denoted by
~$\tilde{\nabla}^{(j)}_{t,i}=\frac{\nabla^{(j)}_{t,i}}{\mathbb{P}[i\in O^{(j)}_{t}]}\cdot\mathbb{I}_{i\in O^{(j)}_{t}}$
for all $j\in[M],i\in[K]$. Clients send $\{\nabla^{(j)}_{t,i},i\in O^{(j)}_{t}\},j\in[M]$ to server. Then server aggregates {~$\{\tilde{\nabla}^{(j)}_{t,i},i\in[K]\}$, $j\in[M]$ and updates the hypotheses following (5)-(7). For each $i\in[K]$, let $\Omega=\mathcal{F}_{i}$ and $\tilde{g}^{(j)}_{t}=\tilde{\nabla}^{(j)}_{t,i}$. Server executes (12).
{¯$\left\{\begin{split}\bar{\nabla}_{t,i}=&\frac{1}{M}\sum^{M}_{j=1}\tilde{\nabla}^{(j)}_{t,i},\\ \nabla_{\bar{{\bf w}}_{t+1,i}}\psi_{t,i}(\bar{{\bf w}}_{t+1,i})=&\nabla_{{\bf w}_{t,i}}\psi_{t,i}({\bf w}_{t,i})-\bar{\nabla}_{t,i},\\ {\bf w}_{t+1,i}=&\mathop{\arg\min}_{{\bf w}\in\mathcal{F}_{i}}\mathcal{D}_{\psi_{t,i}}({\bf w},\bar{{\bf w}}_{t+1,i}),\\ \psi_{t,i}({\bf w})=&\frac{1}{2\lambda_{t,i}}\cdot\|{\bf w}\|^{2}_{\mathcal{F}_{i}},\end{split}\right.$ (12)
where $\psi_{t,i}({\bf w})=\frac{1}{2\lambda_{t,i}}\|{\bf w}\|^{2}_{2}$ is the Euclidean regularizer and $\lambda_{t,i}$ is a time-variant learning rate.
We name this algorithm FOMD-OMS (FOMD-No-LU for OMS-DecD) and show it in Algorithm 3.
Algorithm 3 FOMD-OMS $(R=T)$
0:  $T$, $J$, $\eta_{1}$, $\{U_{i},\lambda_{1,i},i\in[K]\}$
0:  $f^{(j)}_{1,i}=0$, $p_{1,i}$, $i\in[K]$, $j\in[M]$
1:  for $t=1,2,\ldots,T$ do
2:     for $j=1,\ldots,M$ do
3:        Server samples $O^{(j)}_{t}$ following (10)
4:        Server broadcasts $f^{(j)}_{t,i},i\in O^{(j)}_{t}$ to the $j$-th client
5:     end for
6:     for $j=1,\ldots,M$ in parallel do
7:        The client outputs $f^{(j)}_{t,A_{t,1}}({\bf x}^{(j)}_{t})$
8:        The client computes and transmits $\{\nabla^{(j)}_{t,i},c^{(j)}_{t,i}\}_{i\in O^{(j)}_{t}}$
9:     end for
10:     Server computes ${\bf p}_{t+1}$ following (11)
11:     Server computes ${\bf w}_{t+1,i},i\in[K]$ following (12)
12:  end for
5.3 Regret bounds
To obtain high-probability regret bounds that adapt to the complexity of individual hypothesis space, we establish a new Bernstein’s inequality for martingale.
Lemma 1.
Let $X_{1},\ldots,X_{n}$ be a bounded martingale difference sequence w.r.t. the filtration $\mathcal{H}=(\mathcal{H}_{k})_{1\leq k\leq n}$ and with $|X_{k}|\leq a$. Let $Z_{t}=\sum^{t}_{k=1}X_{k}$ be the associated martingale. Denote the sum of the conditional variances by $\Sigma^{2}_{n}=\sum^{n}_{k=1}\mathbb{E}\left[X^{2}_{k}|\mathcal{H}_{k-1}\right]\leq v,$ where $v\in[0,B]$ is a random variable and $B\geq 2$ is a constant. Then for any constant $a>0$, with probability at least $1-2\lceil\log{B}\rceil\delta$,
$\max_{t=1,\ldots,n}Z_{t}<\frac{2a}{3}\ln\frac{1}{\delta}+\sqrt{\frac{2}{B}\ln\frac{1}{\delta}}+2\sqrt{v\ln\frac{1}{\delta}}.$
Note that $v$ is a random variable in Lemma 1, while it is a constant in standard Bernstein’s inequality for martingale (see Lemma A.8 [38]). Lemma 1 is derived from the standard Bernstein’s inequality along with the well-known peeling technique [39].
Assumption 1.
For each $i\in[K]$, there is a constant $b_{i}$ such that $\|\phi_{i}({\bf x})\|_{2}\leq b_{i}$ where $\phi_{i}(\cdot)$ is defined in (1).
Lemma 2.
Under Assumption 1, for each $i\in[K]$, there are two constants $C_{i}>0,G_{i}>0$ that depend on $U_{i}$ or $b_{i}$ such that $\max_{t,j}c^{(j)}_{t,i}\leq C_{i}$ and $\max_{t,j}\|\nabla^{(j)}_{t,i}\|_{2}\leq G_{i}$.
Theorem 3.
Let $\ell(\cdot,\cdot)$ be convex. Under Assumption 1, denote by $A_{m}=\mathrm{argmin}_{i\in[K]}C_{i}$. Let ${\bf p}_{1}$ satisfy
$p_{1,i}=\frac{1-\frac{\sqrt{K}}{\sqrt{T}}}{|A_{m}|}+\frac{1}{\sqrt{KT}},i\in A_{m},p_{1,i}=\frac{1}{\sqrt{KT}},i\neq A_{m}.$
Let $K\geq J\geq 2$ and
$\forall t\in[T],~{}\eta_{t}=$ $\frac{\sqrt{\ln{(KT)}}}{2\sqrt{\left(1+\frac{K-J}{(J-1)M}\right)T}}\wedge\frac{J-1}{2(K-J)},$
$\lambda_{t,i}=$ $\frac{U_{i}}{2G_{i}\sqrt{\left(1+\frac{K-J}{(J-1)M}\right)\cdot\left(\frac{(K-J)^{2}}{(J-1)^{2}}\vee t\right)}}.$
With probability at least $1-\Theta\left(M\log(T)+\log(KT/M)\right)\cdot\delta$, the regret of FOMD-OMS ($R=T$) satisfies: $\forall i\in[K]$,
$\mathrm{Reg}_{D}$ $(\mathcal{F}_{i})=O\left(MB_{i,1}\sqrt{\left(1+\frac{K-J}{(J-1)M}\right)T}+\frac{B_{i,2}(K-J)}{J-1}\ln\frac{1}{\delta}+B_{i,3}\sqrt{\frac{(K-J)MT}{J-1}\ln\frac{1}{\delta}}\right),$
where $B_{i,1}=U_{i}G_{i}+C_{i}\sqrt{\ln(KT)}$, $B_{i,2}=MC+U_{i}G_{i}$, $B_{i,3}=U_{i}G_{i}+\sqrt{CC_{i}}$ and $C=\max_{i\in[K]}C_{i}$.
Both $C_{i}$ and $G_{i}$ depend on $U_{i}$ or $b_{i}$ (see Lemma 2). Let $\mathfrak{C}_{i}=\Theta(U_{i}G_{i}+C_{i})$. Thus $\mathfrak{C}_{i}$ measures the complexity of $\mathcal{F}_{i}$. Then our regret bound adapts to $\sqrt{\mathfrak{C}\mathfrak{C}_{i}}$ where $\mathfrak{C}=\max_{i\in[K]}\mathfrak{C}_{i}$, while previous regret bounds depend on $\mathfrak{C}$ [24, 23] that is, they can not adapt to the complexity of individual hypothesis space. If $\mathfrak{C}_{i^{\ast}}\ll\mathfrak{C}$, then our regret bound is much better.
The regret bound in Theorem 3 is also called multi-scale regret bound [37]. However, previous regret analysis can not yield a high-probability multi-scale bound. The reason is the lack of the new Bernstein’s inequality for martingale (Lemma 1). If we use the new Freedman’s inequality for martingale [40], then a high-probability bound can still be obtained, but is worse than the bound in Theorem 3 by a factor of order $O(\mathrm{poly}(\ln{T}))$.
5.4 Complexity Analysis
For each $j\in[M]$, the $j$-th client makes prediction and computes gradients in time $O(\sum_{i\in O^{(j)}_{t}}d_{i})$. Server samples $O^{(j)}_{t},j\in[M]$, aggregates gradients and updates global models. The per-round time complexity on server is $O(\sum^{M}_{j=1}\sum_{i\in O^{(j)}_{t}}d_{i}+\sum^{K}_{i=1}d_{i}+JM\log{K})$.
Upload At any round $t\in[T]$, the $j$-th client transmits $\tilde{c}^{(j)}_{t,i},\tilde{\nabla}^{(j)}_{t,i}$, $i\in O^{(j)}_{t}$ and the corresponding indexes to server. It requires $J(\sum_{i\in O^{(j)}_{t}}d_{i}+1)$ floating-point numbers and $J$ integers. If we use $32$ bits to represent a float, and use $\log{K}$ bits to represent an integer in $[K]$. Each client transmits $(32J(\sum_{i\in O^{(j)}_{t}}d_{i}+1)+J\log{K})$ bits to server.
Download Server broadcasts ${\bf w}_{t,i}\in\mathbb{R}^{d_{i}},i\in O^{(j)}_{t}$ and the corresponding indexes to clients. The total download cost is $(32MJ(\sum_{i\in O^{(j)}_{t}}d_{i}+1)+MJ\log{K})$ bits.
5.5 Answers to Question 1
Before discussing Question 1, we give two lower bounds on the regret.
Theorem 4 (Lower Bounds).
Assuming that $5\leq K\leq\min\{d,T\}$. For each $i\in[K]$, let $\mathcal{F}_{i}=\{f_{i}({\bf x})={\bf e}^{\top}_{i}{\bf x}\}$ and $\mathcal{D}_{i}=[\min_{{\bf x}\in\mathcal{X}}f_{i}({\bf x}),\max_{{\bf x}\in\mathcal{X}}f_{i}({\bf x})]$, where ${\bf e}_{i}$ is the standard basis vector in $\mathbb{R}^{d}$. Denote by $\sup$ the supremum over all examples.
(i) There are no computational constraints on clients. Let $\ell(v,y)=|v-y|$. The regret of any algorithm for OMS-DecD satisfies: $\lim_{T\rightarrow\infty}\sup\max_{i\in[K]}\mathrm{Reg}_{D}(\mathcal{F}_{i})\geq 0.25M\sqrt{T\ln{K}}$;
(ii) The per-round time complexity on each client is limited to $O(J)$. Let $\ell(v,y)=1-v\cdot y$. The regret of any, possibly randomized, noncooperative algorithm with outputs in $\cup_{i\in[K]}\mathcal{D}_{i}$ satisfies: with probability at least $1-\delta$, $\sup\mathbb{E}[\max_{i\in[K]}\mathrm{Reg}_{D}(\mathcal{F}_{i})]\geq 0.1M\sqrt{KTJ^{-1}}+M\sqrt{0.5T\ln{(M/\delta)}}$, where the expectation is taken over the randomization of algorithm.
The assumption that the outputs of any noncooperative algorithm belong to $\cup_{i=\in[K]}\mathcal{D}_{i}$ is natural, and can be removed in the case of $J=1$. Next we define a noncooperative algorithm, NCO-OMS.
Definition 1 (NCO-OMS).
NCO-OMS independently samples $O^{(j)}_{t}$ following (10) and executes
$\forall j\in[M],\quad\nabla_{\bar{{\bf p}}_{t+1}}\psi_{t}(\bar{{\bf p}}_{t+1})=$ $\nabla_{{\bf p}^{(j)}_{t}}\psi_{t}\left({\bf p}^{(j)}_{t}\right)-\tilde{{\bf c}}^{(j)}_{t},\quad\qquad{\bf p}^{(j)}_{t+1}=\mathop{\arg\min}_{{\bf p}\in\Delta_{K}}\mathcal{D}_{\psi_{t}}({\bf p},\bar{{\bf p}}_{t+1}).$
$\nabla_{\bar{{\bf w}}_{t+1,i}}\psi_{t,i}(\bar{{\bf w}}_{t+1,i})=$ $\nabla_{{\bf w}^{(j)}_{t,i}}\psi_{t,i}\left({\bf w}^{(j)}_{t,i}\right)-\tilde{\nabla}^{(j)}_{t,i},\quad{\bf w}^{(j)}_{t+1,i}=\mathop{\arg\min}_{{\bf w}\in\mathcal{F}_{i}}\mathcal{D}_{\psi_{t,i}}({\bf w},\bar{{\bf w}}_{t+1,i}),$
where the definitions of $\tilde{{\bf c}}^{(j)}_{t}$ and ~$\tilde{\nabla}^{(j)}_{t,i}$ follow FOMD-OMS.
Algorithm 4 NCO-OMS
0:  $T$, $J$, $\eta_{1}$, $\{U_{i},\lambda_{1,i},i\in[K]\}$
0:  $f^{(j)}_{1,i}=0$, $p_{1,i}$, $i\in[K]$, $j\in[M]$
1:  for $t=1,2,\ldots,T$ do
2:     for $j=1,\ldots,M$ do
3:        The client samples $O^{(j)}_{t}$ following (10)
4:        The client outputs $f^{(j)}_{t,A_{t,1}}({\bf x}^{(j)}_{t})$
5:        The client computes $f^{(j)}_{t,A_{t,a}}({\bf x}^{(j)}_{t})$ for all $a=2,\ldots,J$
6:        The client computes ~$\tilde{\nabla}^{(j)}_{t,i}$ and $\tilde{c}^{(j)}_{t,i}$ for all $i\in O^{(j)}_{t}$
7:        The client computes ${\bf p}^{(j)}_{t+1}$ and ${\bf w}^{(j)}_{t+1,i},i\in[K]$ following Definition 1
8:     end for
9:  end for
The pseudo-code of NCO-OMS is shown in Algorithm 4. It is easy to prove the regret of NCO-OMS satisfies: with probability at least $1-\Theta\left(M\log(KT)\right)\cdot\delta$,
$\forall i\in[K],~{}\mathrm{Reg}_{D}$ $(\mathcal{F}_{i})=O\left(M\left(B_{i,1}\sqrt{\left(1+g_{K,J}\right)T}+B_{i,2}g_{K,J}\ln\frac{1}{\delta}+B_{i,3}\sqrt{g_{K,J}T\ln\frac{1}{\delta}}\right)\right),$
where $B_{i,1}=U_{i}G_{i}+C_{i}\sqrt{\ln(KT)}$, $B_{i,2}=C+U_{i}G_{i}$ and $B_{i,3}=U_{i}G_{i}+\sqrt{CC_{i}}$. We leave the pseudo-code of NCO-OMS and the corresponding regret analysis in appendix.
Next we discuss Question 1 by considering two cases.
Case 1: There are no computational constraints on clients. Collaboration is unnecessary.
Let $J=\Theta(K)$ in FOMD-OMS and NCO-OMS. By Theorem 3, both FOMD-OMS and NCO-OMS enjoy a $O(MU_{i}G_{i}\sqrt{T}+MC_{i}\sqrt{T\ln(KT)})$ regret. By Theorem 4, FOMD-OMS and NCO-OMS are nearly optimal in terms of the dependence on $M$ and $T$. Thus collaboration is unnecessary.
Case 2: The per-round time complexity on each client is limited to $o(K)$. Collaboration is necessary.
Let $J=o(K)$ in FOMD-OMS and Theorem 4. By Theorem 3, FOMD-OMS enjoys a $O(MB_{i,1}\sqrt{T}+B_{i,3}\sqrt{MKTJ^{-1}\ln{\delta^{-1}}})$ regret, which is smaller than the lower bound on the regret of any noncooperative algorithm (see Theorem 4). Thus collaboration is necessary.
6 OMS-DecD with Communication Constraint
Let $R<T$. The clients communicate with server every $N$ rounds. For any $r\in[R]$, the clients transmit $\{\frac{1}{N}\sum_{t\in T_{r}}\nabla^{(j)}_{t,i},\frac{1}{N}\sum_{t\in T_{r}}c^{(j)}_{t,i}\}_{i\in O^{(j)}_{t}}$ to server at the last round in $T_{r}$. Then the server updates sampling probabilities and hypotheses. We give the pseudo-code Algorithm 5.
Algorithm 5 FOMD-OMS $(R<T)$
0:  $U$, $T$, $R$, $J$.
0:  $f^{(j)}_{1,i}=0$, $p_{1,i}$, $i\in[K]$, $j\in[M]$
1:  for $r=1,2,\ldots,R$ do
2:     for $t\in T_{r}$ do
3:        if $t==(r-1)N+1$ then
4:           for $j=1,\ldots,M$ do
5:              Server samples $O^{(j)}_{t}$ following (10)
6:              Server transmits $f^{(j)}_{t,i},i\in O^{(j)}_{t}$ to the $j$-th client
7:           end for
8:        end if
9:        for $j=1,\ldots,M$ in parallel do
10:           Output $f^{(j)}_{t,A_{t,1}}({\bf x}^{(j)}_{t})$
11:           for $i\in O^{(j)}_{t}$ do
12:              Computing $\nabla^{(j)}_{t,i}$ and $c^{(j)}_{t,i}$
13:           end for
14:           if $t==rN$ then
15:              Communicate to server: $\{\frac{1}{N}\sum_{t\in T_{r}}\nabla^{(j)}_{t,i},\frac{1}{N}\sum_{t\in T_{r}}c^{(j)}_{t,i}\}_{i\in O^{(j)}_{t}}$
16:           end if
17:        end for
18:        if $t==rN$ then
19:           Server computes ${\bf p}_{t+1}$ following (11)
20:           Server computes ${\bf w}_{t+1,i},i\in[K]$ following (12)
21:        end if
22:     end for
23:  end for
Theorem 5.
For any $r\in[R]$, let ${\bf p}_{1}$, $\eta_{r}$ and $\lambda_{r,i}$ follow Theorem 3, in which we replace $T$ with $R$. Under the condition of Theorem 3, with probability at least $1-\Theta\left(\frac{T}{R}M\log(R)+\frac{T}{R}\log(KR/M)\right)\cdot\delta$, the regret of FOMD-OMS $(R<T)$ satisfies
$\mathrm{Reg}_{D}(\mathcal{F})=O\left(MB_{i,1}\sqrt{\left(1+\frac{K-J}{(J-1)M}\right)}\cdot\frac{T}{\sqrt{R}}+\frac{T}{R}\cdot\frac{B_{i,2}M(K-J)}{J-1}\ln\frac{1}{\delta}+\frac{B_{i,3}T}{\sqrt{R}}\sqrt{\frac{M(K-J)}{J-1}\ln\frac{1}{\delta}}\right).$
The regret bound depends on $O(\frac{1}{\sqrt{R}})$. Thus FOMD-OMS explicitly balances the prediction performance and the communication cost.
7 Application to Distributed OMKL
For each $i\in[K]$, let $\mathcal{F}_{i}$ be a RKHS. FOMD-OMS ($R\leq T$) can solve distributed OMKL [24].
Theorem 6.
Let $\{\mathcal{F}_{i}\}^{K}_{i=1}$ be RHKSs. With probability at least $1-\Theta\left(TM\log(R)+T\log(KR/M)\right)\cdot\delta$, FOMD-OMS satisfies, $\forall i\in[K]$,
$\mathrm{Reg}_{D}(\mathcal{F}_{i})=\tilde{O}\left(MB_{i,1}\sqrt{1+\frac{K-J}{(J-1)M}}\cdot\frac{T}{\sqrt{R}}+\frac{B_{i,2}M(K-J)}{R(J-1)/T}+\frac{B_{i,3}T}{\sqrt{R}}\sqrt{\frac{M(K-J)}{J-1}}+\frac{U_{i}G_{i}MT}{\sqrt{D}}\right),$
where $\tilde{O}(\cdot)$ omits $O(\mathrm{poly}(\ln\frac{1}{\delta}))$ and $D=d_{i}$ follows (1).
We defer the algorithm in appendix. Let $R=T$ and $J=2$. We compare FOMD-OMS with eM-KOFL [23] and POF-MKL [24]. Table 3 gives the results.
We observe that FOMD-OMS significantly improves the computational complexity of eM-KOFL and POF-MKL (by a factor of $O(K)$). The per-round time complexity of the two algorithms is $O(DK)$. Recalling the answer to Question 1 (see Section 5.5), collaboration in eM-KOFL and POF-MKL is unnecessary.
Next we compare the regret bound of the three algorithms. Recalling that $\mathfrak{C}_{i}\leq\mathfrak{C}$. The regret bounds of eM-KOFL and POF-MKL can not adapt to the complexity of individual hypothesis space. (i) The regret bound of FOMD-OMS is better than that of POF-MKL in relation to its dependence on $M$ and $\mathfrak{C}_{i}$. (ii) In the case of $K=O\left(\frac{\mathfrak{C}}{\mathfrak{C}_{i}}M\cdot\ln{K}\right)$, the regret bound of FOMD-OMS is better than that of eM-KOFL. (iii) In the case of $K=\Omega\left(\frac{\mathfrak{C}}{\mathfrak{C}_{i}}M\cdot\ln{K}\right)$, the regret bound of FOMD-OMS is worse than that of eM-KOFL. If $K$ is sufficiently large, the regret bound of eM-KOFL is better than that of FOMD-OMS by a factor of $O(\sqrt{K})$.
Table 3: Comparison with previous algorithms. $D$ is the number of random features [26]. $\tilde{O}(\cdot)$ hides $O(\mathrm{poly}(\ln\frac{T}{\delta}))$. Time (s) is the per-round time complexity on client.
Algorithm Regret bound Time (s) Download (bits)
eM-KOFL $\tilde{O}\left(\mathfrak{C}M\sqrt{T\ln{K}}+\frac{\mathfrak{C}_{i}MT}{\sqrt{D}}\right)$ $O(DK)$ $O(DM\log{K})$
POF-MKL $\tilde{O}\left(\mathfrak{C}M\sqrt{KT}+\frac{\mathfrak{C}_{i}MT}{\sqrt{D}}\right)$ $O(DK)$ $O(DMK)$
FOMD-OMS $\tilde{O}\left(\mathfrak{C}_{i}M\sqrt{T\ln{K}}+\sqrt{\mathfrak{C}\mathfrak{C}_{i}MKT}+\frac{\mathfrak{C}_{i}MT}{\sqrt{D}}\right)$ $O(D)$ $O(DM\log{K})$
8 Experiments
In this section, we aim to verify the following three goals which are our main results.
  1. G1
    Collaboration is unnecessary if we allow the computational cost on each client to be $O(K)$.
    For FOMD-OMS with $R=T$, we set $J=K$. In this case, the per-round running time on each client is $O(K)$. We aim to verify that FOMD-OMS enjoys similar prediction performance with the noncooperative algorithm, NCO-OMS (see Definition 1).
  2. G2
    Collaboration is necessary if we limit the computational cost on each client to $o(K)$.
    For FOMD-OMS with $R=T$, we set $J=2$. In this case, the per-round running time on each client is $O(1)$. We aim to verify that FOMD-OMS enjoys better prediction performance than NCO-OMS.
  3. G3
    FOMD-OMS $(R=T)$ improves the regret bounds of algorithms for distributed OMKL.
    FOMD-OMS $(R=T)$ with $J=2$ enjoys similar prediction performance with eM-KOFL [23], and enjoys better prediction performance than POF-MKL [24] at a smaller computational cost on each client.
    Although there are more baseline algorithms, such as vM-KOFL [23], pM-KOFL [23] and OFSKL [24], we do not compare with the three algorithms since they do not perform as well as eM-KOFL and POF-MKL.
8.1 Experimental setting
We will execute three experiments and each one verifies a goal. For simplicity, we do not measure the actual communication cost and use serial implementation to simulate the distributed implementation.
To verify G1 and G2, we use the instance of online model selection given in Example 1. The first experiment verifies G1. We construct $10$ nested hypothesis spaces (i.e., $K=10$) as follows
$\forall i\in[10],\quad\mathcal{F}_{i}=\left\{f({\bf x})=\langle{\bf w},{\bf x}\rangle,\|{\bf w}\|_{2}\leq U_{i}\right\},$
where $U_{i}=\frac{i}{10}$. We use FOMD-OMS with $R=T$ and set $J=K$. Since $J=K$, we have $O^{(j)}_{t}=[K]$ and $\mathbb{P}\left[i\in O^{(j)}_{t}\right]=1$. The learning rates $\eta_{t},\lambda_{t,i},i\in[K]$ of FOMD-OMS follow Theorem 3. For NCO-OMS, we set $J=K$ and set the learning rate $\eta_{t},\lambda_{t,i},i\in[K]$ following Theorem 3 in which $M=1$, i.e.,
$\forall t\in[T],~{}\eta_{t}=\frac{\sqrt{\ln{(KT)}}}{2\sqrt{T}},\quad\lambda_{t,i}=\frac{U_{i}}{2G_{i}\sqrt{t}}.$
We use the square loss function $\ell(f({\bf x}),y)=(f({\bf x})-y)^{2}$. For both FOMD-OMS and NCO-OMS, we tune $G_{i}=(U_{i}+1)\times\{1,2,4,6,8,10\}$ and set $C_{i}=(U_{i}+1)^{2}$.
The second experiment verifies G2. We use FOMD-OMS with $R=T$ and set $J=2$. The learning rates of FOMD-OMS also follow Theorem 3. For NCO-OMS, we also set $J=2$ and set the learning rate $\eta_{t},\lambda_{t,i},i\in[K]$ following Theorem 3 in which $M=1$, i.e.,
$\forall t\in[T],\quad\eta_{t}=\frac{\sqrt{\ln{(KT)}}}{2\sqrt{\left(K-1\right)T}}\wedge\frac{1}{2(K-2)},\quad\lambda_{t,i}=\frac{U_{i}}{2G_{i}\sqrt{\left(K-1\right)\cdot\left((K-2)^{2}\vee t\right)}}.$
Similar to the first experiment, we tune $G_{i}=(U_{i}+1)\times\{1,2,4,6,8,10\}$ and set $C_{i}=(U_{i}+1)^{2}$.
The third experiment verifies G3. We consider online kernel selection (as known as online multi-kernel learning) which is an instance of online model selection given in Example 2. We select the Gaussian kernel with $8$ different kernel widths (i.e., $K=8$),
$\forall i\in[8],\quad\kappa_{i}({\bf x},{\bf v})=\exp\left(-\frac{\|{\bf x}-{\bf v}\|^{2}_{2}}{2\sigma^{2}_{i}}\right),\sigma_{i}=2^{i-2},$
and construct the corresponding hypothesis space $\mathcal{F}_{i}$ and $\mathbb{H}_{i}$ following (21) in which we set $U_{i}=U$ and $D_{i}=D$ for all $i\in[K]$ and tune $U\in\{1,2,4\}$. Note that $U_{i}$ is same for all $i\in[K]$. We replace the initial distribution ${\bf p}_{1}$ in Theorem 3 with a uniform distribution $(\frac{1}{K},\ldots,\frac{1}{K})$. We set $D=100$ for FOMD-OMS, eM-KOFL and POF-MKL. $D$ is the number of random features. We set $J=2$ and $C=U+1$ in FOMD-OMS. Thus the per-round time complexity on each client is $O(D)$ and the per-round communication cost is $O(MD+M\log{K})$. There are three hyper-parameters in eM-KOFL, i.e., $\eta_{g}$, $\eta_{l}$ and $\lambda$. $\eta_{g}$ is the global learning rate, $\eta_{l}$ is the local learning rate and $\lambda$ is a regularization parameter. There are $32M+$ 3 hyper-parameters in POF-MKL, i.e., $\eta_{g}$, $\eta_{j},\xi_{j},j\in[M]$, $m$, $\lambda$ in which $M/m$ plays the same role with $J$ in FOMD-OMS. $\eta_{g}$ is the global learning rate, $\eta_{j}$ is the local learning rate, $\xi_{j}$ is called exploration rate and $\lambda$ is a regularization parameter. Since $J=2$ in FOMD-OMS, we can set $m=M/2$ for FOMD-OMS. Following the original paper [24], we set $\xi_{j}=1$. For a fair comparison, we change the learning rates of FOMD-OMS, eM-KOFL and POF-MKL. Following the parameter setting of eM-KOFL [23], we tune $\eta_{g},\eta_{l},\eta_{j}\in\{0.1,0.5,1,4,8,16\}$ and $\lambda\in\{0.1,0.001,0.0001\}$ for eM-KOFL and POF-MKL. For FOMD-OMS, we also tune $\eta_{t},\lambda_{t,i}\in\{0.1,0.5,1,4,8,16\}$.
For all of the three experiments, we set $10$ clients, i.e., $M=10$. We use 8 regression datasets shown in Table 4 from WEKA and UCI machine learning repository  [^1] https://archive.ics.uci.edu/ml/index.php, and rescale the target variables and features of all datasets to fit in [0,1] and [-1,1] respectively. For each dataset, we randomly divide it into $10$ subsets and each subset simulates the data on a client. We randomly permutate the instances in the datasets 10 times and report the average results. All algorithms are implemented with R on a Windows machine with 2.8 GHz Core(TM) i7-1165G7 CPU.
Table 4: Basic information of datasets.
  Dataset Number of Instances Number of features
bank 8,190 32
ailerons 13,750 40
calhousing 14,000 8
elevators 16,590 18
TomsHardware 28,170 96
Twitter 50,000 77
Year 51,630 90
Slice 53,500 384
We use the square loss function and define the mean squared error (MSE) of all algorithms, i.e.,
$\mathrm{MSE}=\frac{1}{MT}\sum^{M}_{j=1}\sum^{T}_{t=1}\left(f^{(j)}_{t}\left({\bf x}^{(j)}_{t}\right)-y^{(j)}_{t}\right)^{2}.$
We record the mean of MSE over 10 random experiments, and the standard deviation of the mean of MSE. We also record the mean of of the total running time on each client, and the standard deviation of the mean of running time.
8.2 Results of the First and the Second Experiment
We summary the experimental results of the first and the second experiments in Table 5.
Table 5: Comparison with the noncooperative algorithm. $\Delta$ is the difference of MSE between NCO-OMS and FOMD-OMS. $3\mathrm{E}\text{-}4=3\times 10^{-4}$ and $1\mathrm{E}\text{-}5=1\times 10^{-5}$.
  Algorithm elevator bank
MSE×$\times 10^{2}$ $J$ Time (s) $\Delta$ MSE×$\times 10^{2}$ $J$ Time (s) $\Delta$
NCO-OMS 0.991 $\pm$ 0.002 $K$ 1.31 $\pm$ 0.10 0.0001 2.158 $\pm$ 0.022 $K$ 0.88 $\pm$ 0.05 0.001
FOMD-OMS 0.980 $\pm$ 0.005 $K$ 0.65 $\pm$ 0.08 2.020 $\pm$ 0.005 $K$ 0.33 $\pm$ 0.08
NCO-OMS 1.168 $\pm$ 0.005 $2$ 0.58 $\pm$ 0.04 0.001 2.321 $\pm$ 0.021 $2$ 0.39 $\pm$ 0.05 0.002
FOMD-OMS 1.024 $\pm$ 0.002 $2$ 0.14 $\pm$ 0.04 2.118 $\pm$ 0.003 $2$ 0.08 $\pm$ 0.03
  Algorithm TomsHardware Twitter
MSE×$\times 10^{2}$ $J$ Time (s) $\Delta$ MSE×$\times 10^{2}$ $J$ Time (s) $\Delta$
NCO-OMS 0.090 $\pm$ 0.004 $K$ 3.02 $\pm$ 0.29 0.0001 0.017 $\pm$ 0.000 $K$ 5.11 $\pm$ 0.24 0
FOMD-OMS 0.083 $\pm$ 0.008 $K$ 1.48 $\pm$ 0.28 0.017 $\pm$ 0.000 $K$ 2.07 $\pm$ 0.07
NCO-OMS 0.150 $\pm$ 0.002 $2$ 1.11 $\pm$ 0.07 0.0004 0.018 $\pm$ 0.000 $K$ 2.24 $\pm$ 0.25 1E-5
FOMD-OMS 0.107 $\pm$ 0.003 $2$ 0.44 $\pm$ 0.09 0.017 $\pm$ 0.000 $2$ 0.51 $\pm$ 0.05
  Algorithm ailerons calhousing
MSE×$\times 10^{2}$ $J$ Time (s) $\Delta$ MSE×$\times 10^{2}$ $J$ Time (s) $\Delta$
NCO-OMS 19.506 $\pm$ 0.033 $K$ 1.41 $\pm$ 0.04 0.0003 10.166 $\pm$ 0.029 $K$ 1.07 $\pm$ 0.05 3E-4
FOMD-OMS 19.480 $\pm$ 0.046 $K$ 0.74 $\pm$ 0.08 10.136 $\pm$ 0.012 $K$ 0.43 $\pm$ 0.05
NCO-OMS 20.323 $\pm$ 0.036 $2$ 0.65 $\pm$ 0.05 0.0050 10.372 $\pm$ 0.021 $2$ 0.53 $\pm$ 0.04 0.001
FOMD-OMS 19.820 $\pm$ 0.032 $2$ 0.20 $\pm$ 0.04 10.227 $\pm$ 0.014 $2$ 0.12 $\pm$ 0.05
  Algorithm year Slice
MSE×$\times 10^{2}$ $J$ Time (s) $\Delta$ MSE×$\times 10^{2}$ $J$ Time (s) $\Delta$
NCO-OMS 20.322 $\pm$ 0.040 $K$ 5.94 $\pm$ 0.25 0.0023 13.097 $\pm$ 0.009 $K$ 10.40 $\pm$ 0.94 0.001
FOMD-OMS 20.096 $\pm$ 0.045 $K$ 2.25 $\pm$ 0.20 12.964 $\pm$ 0.007 $K$ 4.12 $\pm$ 0.18
NCO-OMS 24.334 $\pm$ 0.021 $2$ 2.95 $\pm$ 0.63 0.0163 13.364 $\pm$ 0.012 $2$ 3.60 $\pm$ 0.23 0.003
FOMD-OMS 22.705 $\pm$ 0.040 $2$ 0.59 $\pm$ 0.09 13.038 $\pm$ 0.009 $2$ 1.41 $\pm$ 0.12
In Table 5, $\Delta$ is defined as the difference of MSE between NCO-OM and FOMD-OMS. Thus $\Delta$ shows whether collaboration improves the prediction performance of the noncooperative algorithm. Times (s) records the total running time on all clients.
We first consider the case $J=K$ in which the per-round time complexity on each client is $O(K)$. It is obvious that the MSE of NCO-OMS is similar with that of FOMD-OMS. Although there are four datasets on which FOMD-OMS performs better than NCO-OMS, such as the elevator, bank, Year and Slice datsets, the improvement is very limited. Beside, the value of $\Delta$ is very small. Thus collaboration does not significantly improve the prediction performance of the noncooperative algorithm. The results verify the first goal G1.
Next we consider the case $J=2$ in which the per-round time complexity on each client is $O(1)$. It is obvious that FOMD-OMS performs better than NCO-OMS on all datasets. Besides, the value of $\Delta$ in the case of $J=2$ is much larger than that in the case of $J=K$, such as the elevators, ailerons, ailerons and Year datasets. Thus collaboration indeed improves the prediction performance of the noncooperative algorithm. The results verify the second goal G2.
Finally we compare the running time of all algorithms. It is obvious that FOMD-OMS with $J=2$ runs faster than the other algorithms. The results coincide with our theoretical analysis. NCO-OMS runs slower than FOMD-OMS. The reason is that NCO-OMS must solve the sampling probability ${\bf p}_{t}$ using an additional binary search on each client (see Section 15.1). In other words, NCO-OMS must execute binary search $M$ times at each round. FOMD-OMS only executes one binary search on server at each round. The improvement on the computational cost is benefit from decoupling model selection and prediction.
8.3 Results of the Third Experiment
We summary the experimental results of the third experiment in Table 6.
Table 6: Comparison with the state-of-the-art algorithms.
  Algorithm elevator bank
MSE $J$ Time (s) MSE $J$ Time (s)
eM-KOFL 0.00292 $\pm$ 0.00013 - 2.67 $\pm$ 0.05 0.01942 $\pm$ 0.00066 - 1.41 $\pm$ 0.06
POF-MKL 0.00806 $\pm$ 0.00026 - 3.12 $\pm$ 0.14 0.02292 $\pm$ 0.00036 - 1.59 $\pm$ 0.13
FOMD-OMS 0.00318 $\pm$ 0.00021 $2$ 0.52 $\pm$ 0.08 0.01917 $\pm$ 0.00110 $2$ 0.27 $\pm$ 0.06
  Algorithm TomsHardware Twitter
MSE $J$ Time (s) MSE $J$ Time (s)
eM-KOFL 0.00048 $\pm$ 0.00003 - 5.88 $\pm$ 0.69 0.00007 $\pm$ $0$ 9.60 $\pm$ 0.77
POF-MKL 0.00188 $\pm$ 0.00004 - 6.60 $\pm$ 0.93 0.00020 $\pm$ 0.00001 $2$ 10.44 $\pm$ 0.54
FOMD-OMS 0.00059 $\pm$ 0.00003 $2$ 1.46 $\pm$ 0.12 0.00010 $\pm$ 0.00001 $2$ 2.23 $\pm$ 0.18
  Algorithm ailerons calhousing
MSE $J$ Time (s) MSE $J$ Time (s)
eM-KOFL 0.00370 $\pm$ 0.00011 - 2.40 $\pm$ 0.19 0.02242 $\pm$ 0.00043 - 2.28 $\pm$ 0.06
POF-MKL 0.01335 $\pm$ 0.00046 - 2.66 $\pm$ 0.12 0.05248 $\pm$ 0.00197 - 2.68 $\pm$ 0.08
FOMD-OMS 0.00429 $\pm$ 0.00021 $2$ 0.48 $\pm$ 0.04 0.02373 $\pm$ 0.00126 $2$ 0.39 $\pm$ 0.07
  Algorithm year Slice
MSE $J$ Time (s) MSE $J$ Time (s)
eM-KOFL 0.01481 $\pm$ 0.00108 - 9.60 $\pm$ 0.51 0.05781 $\pm$ 0.00230 - 12.74 $\pm$ 0.95
POF-MKL 0.01896 $\pm$ 0.00036 - 10.73 $\pm$ 0.29 0.08675 $\pm$ 0.00402 - 14.22 $\pm$ 0.54
FOMD-OMS 0.01534 $\pm$ 0.00121 $2$ 2.26 $\pm$ 0.10 0.05698 $\pm$ 0.00480 $2$ 4.82 $\pm$ 0.21
We first compare FOMD-OMS with eM-KOFL. As a whole, the MSE of the two algorithms is similar. On the TomsHardware, Twitter and ailerons datasets, eM-KOFL enjoys slightly better prediction performance than FOMD-OMS. However, the running time of eM-KOFL is much larger than that of FOMD-OMS. The results coincide with the theoretical observations that FOMD-OMS enjoys a similar regret bound with eM-KOFL at a much smaller computational cost on the clients.
Next we compare FOMD-OMS with POF-MKL. Both the MSE and running time of FOMD-OMS are much smaller than that of POF-MKL. The results coincide with the theoretical observations that FOMD-OMS enjoys a smaller regret bound than POF-MKL at a much smaller computational cost on the clients.
Thus the results in Table 6 verifies the third goal G3.
Finally, we explain that why POF-MKL performs worse than FOMD-OMS. There are three reasons.
  1. (1)
    POF-MKL does not use federated learning to learn a global probability distribution denoted by ${\bf p}_{t}$, but learns a personalized probability distribution denoted by ${\bf p}_{t,j}$ on each client. Thus POF-MKL converges to the best kernel function at a lower rate.
  2. (2)
    POF-MKL uniformly samples two kernel functions and then learns two global hypotheses, while FOMD-OMS uses ${\bf p}_{t}$ to sample a kernel function and learns a global hypothesis. Thus POF-MKL can learn a better global hypothesis.
  3. (3)
    On each client, POF-MKL executes model selection and combines the predictions of $K$ hypotheses using ${\bf p}_{t,j}$. Thus the time complexity is in $O(DK)$. FOMD-OMS executes model selection on server, and only uses the sampled hypothesis to make prediction. Thus the time complexity on each client is in $O(D)$.
9 Conclusion
In this paper, we have studied the necessity of collaboration in OMS-DecD from the perspective of computational constraints. We demonstrate that collaboration is unnecessary when there are no computational constrains on clients, while it becomes necessary if the time complexity on each client is limited to $o(K)$. Our work clarifies the unnecessary nature of collaboration in previous algorithms for the first time, gives conditions under which collaboration is necessary, and provides inspirations for studying the problem from constraints beyond computational constrains.